OpenStudy (anonymous):

Sketch one cycle of the sine curve y = 2 sin theta

OpenStudy (rajee_sam):

y = 2 Sin Θ Sine curve has a period of 2π. So like you graph any other function, give different values for Θ between 0 and 2π and find the value for y For Θ = 0 ; y = 2 Sin 0 = 2 . 0 = 0 For Θ = π/2 ; y = 2 Sin π/2 = 2.1 = 2 For Θ = π ; y = 2 Sin π = 2.0 = 0 For Θ = 3π/2 ; y = 2 Sin 3π/2 = 2. -1 = -2 For Θ = 2π ; y = 2 Sin 2π = 2.0 = 0 Now graph this.